# On the reconstruction map in JT gravity

### v1

@@ A.2 Comparing quantum and classical angle dynamics
Here, we will review some properties of the persistence amplitude $z(t)$ in (4.8), which measures the “closeness” of the quantum dynamics of the angle coherent states to classical action-angle dynamics, from [4]. In generic spectra, the $\Delta_{n}$ defined in (4.9) follow a Gaussian distribution. This is particularly true to a high degree of accuracy for RMT statistics [29], and is also a reasonable approximation for Poisson statistics. In this case, the persistence amplitude behaves as a Gaussian in time:
@@
being a measure of spectral rigidity [30, 31] within the microcanonical window.
@@
which is the precise range of spectral rigidity for RMT statistics [30, 31, 32]. Specifically, for the circular unitary ensemble (CUE), whose eigenvalue statistics is expected for JT gravity, and for Poisson statistics representing a generic non-ergodic system, we have
@@ A.3 Early time dynamics from spectral self-similarity
Both RMT and Poisson spectra show a degree of self-similarity: if one selects any subset of $d_{s}$ consecutive energy levels from such spectra, they continue to show the respective RMT or Poisson behavior in statistical measures such as the distribution of $\Delta_{n}$ and spectral rigidity (with $d$ replaced by $d_{s}$). They also retain the same Heisenberg time $t_{{\rm H}}$ as the full spectrum. We can use this property to quantitatively derive how an initial coherent state $\lvert\tilde{\theta}(0)\rangle$ spreads in $\tilde{\theta}$ as a function of time.
@@
Now, suppose that $20<\kappa\leq 1/$ 2, and that
@@
Figure 7: Log-Log plot of the size $\Delta\tilde{\theta}(t)$ of the smallest symmetric region around $\tilde{\theta}_{\text{cl}}(t)$ in which there is a probability of greater than $0.951-\epsilon=$ 0.95 of measuring $\tilde{\theta}$, for CUE and Poisson statistics, as a function of the normalized time $\tau=td/t_{{\rm H}}$. These are respectively compared with the analytical estimates above by plotting functions of the form Eq. (A.41a) and Eq. (A.41b) with constants $c_{1}$ and $c_{2}$ chosen by hand. These are for a microcanonical window with $d=1024$; see also Fig. 6 for another illustration of the spread.
@@ A.4 A bump at late times
To detect this feature analytically, it is useful to look at a coarse-grained measure of ergodicity (see Appendix B of [33]) by taking the Fourier components of a state $\lvert\tilde{\theta}_{0}(t)\rangle$ along $\tilde{\theta}$. This metric is sensitive to level statistics: for integer $r>0$,
@@
where $S_{r}=E_{n+r}-E_{n}$ denotes the $r$-th nearest neighbor level spacings (numbering $d-r$ in a set of $d$ consecutive levels) and $P^{(r)}(S_{r})$ is their probability distribution [30]. An analogous expression is obtained for $r<0$. In the large $d$ limit, and upon taking $\tilde{\theta}_{0}\rightarrow 0$ as is appropriate for our application, this implies
@@
The observable $\cos[r\tilde{\theta}(t)]$ partitions the circle $\tilde{\theta}\in[0,2\pi)$ into $r$ regions where it is positive and another $r$ regions where it is negative, such that both kinds of regions have the same length. For “coarse-grained” ergodicity, we want $\tilde{\theta}(t)$ to “equally” explore both types of regions over time, which means the expectation value $\langle\cos[r\tilde{\theta}(t)]\rangle_{\tilde{\theta}(0)=0}$ should integrate to zero [^12] It is perhaps more natural to consider a coarse-grained version of ergodicity in terms of Walsh-Hadamard functions (which take the values $+1$ and $-1$ on these same intervals) rather than cosines, but these will involve more complicated combinations of level spacings so we stick with cosines. But our conclusions will turn out to be similar in either case.. Note that we eventually expect $\tilde{\theta}$ to become uniformly distributed as $t\to\infty$ in all generic spectra without degeneracies, so a time average of such observables (as in conventional ergodicity) will always decay to zero; a time integral is therefore a more appropriate quantity to probe “coarse-grained ergodicity”.
@@
$A(\tilde{\theta}):=$ $\int_{-\pi}^{\pi}\frac{\mathrm{d}\tilde{\theta}}{2\pi d^{-1}}\ a(\tilde{\theta})\lvert\tilde{\theta}\rangle\langle\tilde{\theta}\rvert$
@@
Now we specialize to an observable $A(\tilde{\vartheta}(t))$ that depends on the angle variable only through $\tilde{\vartheta}=\lvert\tilde{\theta}\rvert$ (i.e., is invariant under $\tilde{\theta}\to-\tilde{\theta}$) — such as the reconstruction of the wormhole length operator $x(N,\theta)$ in (3.25) (where the $N$-dependence can be dropped in micrcanonical windows, see also Appendix C). If $A(\tilde{\vartheta})$ remains sufficiently smooth as $d\to\infty$, it can be written as a Fourier series that retains only even combinations of the terms in (A.45):
@@
Over these scales, both CUE and Poisson spectra have $P^{(r>1)}(0)=0$, but differ [30] for $r=1$. While a CUE spectrum has complete level repulsion with $P^{(1)}(0)=0$, a Poisson spectrum lacks nearest-neighbor level repulsion and has $P^{(1)}(0)>0$. It follows that every observable $\langle A(\tilde{\vartheta}(t))\rangle$ must oscillate “equally” around its saturation value for a CUE spectrum (such as in JT gravity), indicating “coarse-grained ergodicity”, but this can not happen universally for all such observables given a Poisson spectrum. [^14] We also note that a largely complementary argument in [34], originally for suitably averaged return probabilities over shorter timescales, can be applied to our case to show that if all the $P^{(r)}(S)$ attain a maximum away from $S=0$, irrespective of level repulsion at $S=0$, then $[A(\tilde{\theta}(t))-a_{0}]$ cannot be everywhere positive or everywhere negative in $t$. However, such an argument does not estimate the magnitude of this effect, e.g., whether $[A(\tilde{\theta}(t))-a_{0}]$ explores both sides equally, which would not be the case in the absence of level repulsion. This constitutes a coarse-grained form of ergodic quantum dynamics, in which level repulsion in the spectrum leads to observables “equally” exploring both sides of their long-time saturation value.
For more intuition, let us focus on the observable $\cos[\tilde{\theta}(t)]$ (i.e. $a_{r}=\delta_{r,1}$ in (A.47)), which is $>0$ when $\tilde{\theta}(t)$ is closer to the initial state $\tilde{\theta}=0$, and $<0$ when it is closer to the “diametrically opposite” value $\tilde{\theta}\approx\pi$. If $\langle\cos\tilde{\theta}\rangle<0$, it suggests that the system has a larger probability of $\vartheta>\pi/2$ than $\vartheta<\pi/2$ (although mathematically this is not guaranteed). Using the nearest-neighbor level spacing statistics for CUE and Poisson distributions in the $d\to\infty$ limit:
@@
At $t=0$, this starts of at $1$, but crosses over to negative values at $t=t_{{\rm H}}\sqrt{(2/\pi^{3})}\approx 0.254t_{{\rm H}}$ — just a little higher than the time $t=t_{{\rm H}}/4$ when the classical trajectory would cross $\tilde{\theta}=\pi/4$. Subsequently, it reaches a maximum negative value of $-2\mathrm{e}^{-3/2}\approx-0.446$ at $t\approx 0.440t_{{\rm H}}$, after which it gradually decays to 00. Thus, $\cos(\tilde{\theta})$ captures a coarse grained variant of ergodicity in which the initial wavepacket at $\tilde{\theta}=0$ manages to explore the “opposite side” near $\tilde{\theta}=\pi$ before truly randomizing over the full range of angles. This is illustrated in Fig. 8.
(a) The bump in $\langle\cos(\tilde{\theta})\rangle$.
@@
Figure 8: Illustration of “coarse-grained ergodicity” via the late-time bump feature for the observables $\cos(\tilde{\theta})$ and $\vartheta$ for microcanonical windows of $d=1024$ levels with CUE statistics, and its absence with Poisson statistics, as a function of the normalized time $\tau=td/t_{{\rm H}}$.
@@ Appendix B Non-uniqueness of $R^{*}$ in JT gravity
Happily, if the Fourier coefficients $A_{m}$ and $B_{m}$ decay relatively quickly with $m$, i.e. the functions of $\theta$ are reasonably smooth, and $n\gg 1$, these discrepancies will be negligible in practice. We remind the reader that for the semiclassical states of interest, $n\sim\mathrm{e}^{S}\gg 1$, and the length operator matrix elements (3.36) decay such that $\langle 0|x|n\rangle\sim n^{-2}$, meaning that discrepancies in operator reconstruction will be suppressed by factors of $\mathrm{e}^{-\mathrm{O}(S)}$.
@@ Appendix C Microcanonical approximations to length operators
Let $\Pi_{r}$ project onto a Hilbert space $\mathcal{H}_{r}$ spanned by $d$ consecutive energy levels from $E_{r}$ to $E_{r+d-1}$, each such set of levels constituting a microcanonical window, such that $\sum_{r}\Pi_{r}=1$. The index $r=(m-1)d$ represents the “base index” of the $m$-th microcanonical window in the energy spectrum. In practice, we will be interested in a subset of these microcanonical windows spanning some interval $[E_{\min},E_{\max}]$; let $R$ be the set of values of the base index $r$ corresponding to this interval.
@@
Let $\mathcal{H}_{\text{mc}}(\nu_{\text{mc}})$ be the set of states that are supported on $\nu_{\text{mc}}$ consecutive microcanonical windows spanning $[E_{\min},E_{\max}]$ and are unbiased in each window:
@@
=$=\begin{dcases}\frac{\pi}{2},\ &\text{ for }n=m,\\ -\frac{\pi(1+(-1)^{m-n})}{d^{2}\sin^{2}\left(\frac{m-n}{d}\pi\right)},\ &\text{ for }n\neq m.\end{dcases}$ (C.13)
@@
In the second line, we have set $q=m-n$, used the evenness of the sum in $q$, and noted that $\operatorname{sinc}^{2}(0)=1$ by definition to allow $q=0$. Introducing the variable $\xi=q/d$, we can write the sum in the above expression as:
@@
for some small $\epsilon_{\beta}>0$ (note that the density of states is monotonically increasing with energy, so if this can be established for the $d$ lowest levels in $[E_{\min},E_{\max}]$, we should expect it to be the case for [almost] all energy levels throughout the spectrum, especially given spectral rigidity, as successive sets of $d$ levels span increasingly narrower energy ranges). We get
@@
As we discuss in Sec. 4.2, most of the probability of the Hartle-Hawking state can be found in a finite ($\beta$-dependent) energy range, which we choose to be $[E_{\min},E_{\max}]$, we can truncate the support of $\lvert\beta_{\text{mc}}\rangle$ to this range and normalize this truncated version so that it formally belongs to $\mathcal{H}_{\text{mc}}(\nu_{\mathrm{mc}})$. But it is also the case that to accurately capture expectation values such as $\langle x^{q}\rangle$, the choice of microcanonical windows that must be accounted for is rather different: see Appendix D.
@@ Appendix E Wormhole dynamics with Poisson statistics
At a technical level, we can simulate this system by following the procedure of Sec. 4.3 except with the CUE eigenphases $\varphi_{n}$ replaced by a uniform, random (and sorted) distribution of points in $[0,2\pi)$. With Poisson statistics, we expect the initial state to spread out rapidly to the point of exhibiting no clear imprints of the classical dynamics near the Heisenberg time scale due to a lack of spectral rigidity, and the absence of a bump in $\langle x(t)\rangle$ (more precisely, the absence of an “equal” exploration of values below and above the saturation value) due to a lack of nearest-neighbor level repulsion. This is illustrated in Figs. 9 and 10, respectively for a microcanonical state and a Hartle-Hawking state.

### v2

@@ A.2 Comparing quantum and classical angle dynamics
Here, we will review some properties of the persistence amplitude $z(t)$ in (4.8), which measures the “closeness” of the quantum dynamics of the angle coherent states to classical action-angle dynamics, from [dynamicalqergodicity]. Its behavior may be succinctly summarized as follows. In the case of random matrix statistics, the angle coherent states follow classical dynamics extremely closely up to a time that is almost as large as the Heisenberg time, being smaller only by a low power-law in the entropy, $\exp(S(E))/\sqrt{S(E)}$; this will lead to the “slow wavepacket spread” phenomenon in Sec. A.3. These states also retain some noticeable memory of classical dynamics for at least half the period of the angle variable both forward and backward in time, which is similar (though not identical) to the bump phenomenon to be discussed in Sec. A.4. However, with Poisson statistics, strong deviations from classical dynamics are observed at the exponentially smaller time scale $\exp(S(E)/2)$, thereby being quite far from either of these features.
In generic spectra, the $\Delta_{n}$ defined in (4.9) follow a Gaussian distribution. This is particularly true to a high degree of accuracy for RMT statistics [GRMGaussian], and is also a reasonable approximation for Poisson statistics. In this case, the persistence amplitude behaves as a Gaussian in time:
@@
being a measure of spectral rigidity [Mehta, DeltaStar2] within the microcanonical window.
@@
which is the precise range of spectral rigidity for RMT statistics [Mehta, DeltaStar2, ShenkerThouless]. Specifically, for the circular unitary ensemble (CUE), whose eigenvalue statistics is expected for JT gravity, and for Poisson statistics representing a generic non-ergodic system, we have
@@ A.3 Early time dynamics from spectral self-similarity
Both RMT and Poisson spectra show a degree of self-similarity: if one selects any subset of $d_{s}$ consecutive energy levels from such spectra, they continue to show the respective RMT or Poisson behavior in statistical measures such as the distribution of $\Delta_{n}$ and spectral rigidity (with $d$ replaced by $d_{s}$). They also retain the same Heisenberg time $t_{{\rm H}}$ as the full spectrum. We can use this property to quantitatively derive how an initial coherent state $\lvert\tilde{\theta}(0)\rangle$ spreads in $\tilde{\theta}$ as a function of time. In particular, we will substantiate our claim that the spread of the coherent state is parametrically slower with random matrix statistics (as expected for JT gravity) than for a Poisson spectrum.
@@
Now, suppose that $0<\kappa\leq 1/2$, and that
@@
Figure 7: Log-Log plot of the size $\Delta\tilde{\theta}(t)$ of the smallest symmetric region around $\tilde{\theta}_{\text{cl}}(t)$ in which there is a probability of greater than $1-\epsilon=0.95$ of measuring $\tilde{\theta}$, for CUE and Poisson statistics, as a function of the normalized time $\tau=td/t_{{\rm H}}$. These are respectively compared with the analytical estimates above by plotting functions of the form Eq. (A.41a) and Eq. (A.41b) with constants $c_{1}$ and $c_{2}$ chosen by hand. These are for a microcanonical window with $d=1024$; see also Fig. 6 for another illustration of the spread.
@@ A.4 A bump at late times
Intuitively, this is because the time integral of the deviation of any such expectation value from its long-time saturation value is determined by the density of level spacings in the spectrum near 0, due to the standard Fourier relationship between energy and time. Consequently, if level repulsion (vanishing density of spacings near 0) is present in the spectrum, these deviations must “equally” explore either side of the saturation value so that they cancel each other out to make the time integral vanish. This is directly responsible for the bump feature. In the absence of level repulsion, these deviations need not exactly cancel each other out over time, leading to the absence of a prominent bump.
To detect this feature analytically, it is useful to look at a coarse-grained measure of ergodicity (see Appendix B of [qergthesis]) by taking the Fourier components of a state $\lvert\tilde{\theta}_{0}(t)\rangle$ along $\tilde{\theta}$. This metric is sensitive to level statistics: for integer $r>0$,
@@
where $S_{r}=E_{n+r}-E_{n}$ denotes the $r$-th nearest neighbor level spacings (numbering drd-r in a set of $d$ consecutive levels) and $P^{(r)}(S_{r})$ is their probability distribution [Mehta]. An analogous expression is obtained for $r<0$. In the large $d$ limit, and upon taking $\tilde{\theta}_{0}\rightarrow 0$ as is appropriate for our application, this implies
@@
The observable $\cos[r\tilde{\theta}(t)]$ partitions the circle $\tilde{\theta}\in[0,2\pi)$ into $r$ regions where it is positive and another $r$ regions where it is negative, such that both kinds of regions have the same length. For “coarse-grained” ergodicity, we want $\tilde{\theta}(t)$ to “equally” explore both types of regions over time, which means the expectation value $\langle\cos[r\tilde{\theta}(t)]\rangle_{\tilde{\theta}(0)=0}$ should integrate to zero [^12] It is perhaps more natural to consider a coarse-grained version of ergodicity in terms of Walsh-Hadamard functions (which take the values +1+1 and 1-1 on these same intervals) rather than cosines, but these will involve more complicated combinations of level spacings so we stick with cosines. But our conclusions will turn out to be similar in either case.. Note that we eventually expect $\tilde{\theta}$ to become uniformly distributed as $t\to\infty$ in all generic spectra without degeneracies, so a time average of such observables (as in conventional ergodicity) will always decay to zero; a time integral is therefore a more appropriate quantity to probe “coarse-grained ergodicity”.
@@
[…]
@@
[…]
@@
Over these scales, both CUE and Poisson spectra have $P^{(r>1)}(0)=0$, but differ [Mehta] for $r=1$. While a CUE spectrum has complete level repulsion with $P^{(1)}(0)=0$, a Poisson spectrum lacks nearest-neighbor level repulsion and has $P^{(1)}(0)>0$. It follows that every observable $\langle A(\tilde{\vartheta}(t))\rangle$ must oscillate “equally” around its saturation value for a CUE spectrum (such as in JT gravity), indicating “coarse-grained ergodicity”, but this can not happen universally for all such observables given a Poisson spectrum. [^14] We also note that a largely complementary argument in [WilkieBrumerReturnProbabilities], originally for suitably averaged return probabilities over shorter timescales, can be applied to our case to show that if all the $P^{(r)}(S)$ attain a maximum away from $S=0$, irrespective of level repulsion at $S=0$, then $A(\tilde{\theta}(t))-a_{0}]$ cannot be everywhere positive or everywhere negative in $t$. However, such an argument does not estimate the magnitude of this effect, e.g., whether $A(\tilde{\theta}(t))-a_{0}]$ explores both sides equally, which would not be the case in the absence of level repulsion. This constitutes a coarse-grained form of ergodic quantum dynamics, in which level repulsion in the spectrum leads to observables “equally” exploring both sides of their long-time saturation value.
For more intuition, let us focus on the observable $\cos[\tilde{\theta}(t)]$ (i.e. $a_{r}=\delta_{r,1}$ in (A.47)), which is >0>0 when $\tilde{\theta}(t)$ is closer to the initial state $\tilde{\theta}=0$, and <0<0 when it is closer to the “diametrically opposite” value $\tilde{\theta}\approx\pi$. If $\langle\cos\tilde{\theta}\rangle<0$, it suggests that the system has a larger probability of $\vartheta>\pi/2$ than $\vartheta<\pi/2$ (although mathematically this is not guaranteed). Using the nearest-neighbor level spacing statistics for CUE and Poisson distributions in the $d\to\infty$ limit:
@@
[…]
@@
[…]
@@ Appendix B Non-uniqueness of $R^{*}$ in JT gravity
Happily, if the Fourier coefficients $A_{m}$ and $B_{m}$ decay relatively quickly with $m$, i.e. the functions of $\theta$ are reasonably smooth, and $n\gg 1$, these discrepancies will be negligible in practice when acting on high-energy states, including semiclassical states. We remind the reader that these states have the form:
$\lvert\psi\rangle=\sum_{n}c_{n}\lvert n\rangle,$ (B.4)
where the distribution of $\lvert c_{n}\rvert^{2}$ is peaked (and has most of its weight) at $n\sim\mathrm{e}^{S}\gg 1$. Then, as long as $A_{m}(n)$ and $B_{m}(n)$ are slowly varying with $n$ (which we expect to be the case at large $n$ for most reconstructed observables of physical interest), we have $A_{m_{1}}(n+\delta n)B_{m_{2}}(n+\delta n^{\prime})\approx A_{m_{1}}(n)B_{m_{2}}(n)$, due to which
$R^{*}(A)R^{*}(B)\lvert\psi\rangle\approx R^{*}(AB)\lvert\psi\rangle,$ (B.5)
by (B.3). However, for low-energy states, the ground state cutoffs such as $\delta_{n+m_{2}\geq 0}$ in (B.3) play a more important role, and we are no longer guaranteed that e.g. $A_{m}(n)$ is slowly varying with $n$, making the two actions different in general. As long as we can neglect the ground state truncation of the Hilbert space and have smooth coefficients $A_{m}(n)$, the algebra homomorphism is essentially exact up to operator ordering ambiguities, which are $\sim e^{-\mathrm{O}(S)}$.
For example, the length operator has coefficients $A_{m}(n)$ that are smooth in $n$ but fluctuate only in $m$, and its matrix elements (3.36) decay such that $\langle 0|x|n\rangle\sim n^{-2}$, meaning that discrepancies in operator reconstruction will be suppressed by factors of $\mathrm{e}^{-\mathrm{O}(S)}$ relative to the expectation value of the operator.
Finally, the approximate algebra homomorphism also holds at late times; the late-time deviations from semiclassical dynamics are entirely due to the discrete fundamental Hilbert space and its implications for the expectation values of observables, including $A$ and $B$ separately or either reconstruction of the product $AB$. It is not clear to us that this approximate homomorphism would hold for the prescription chosen in [Iliesiu:2021ari, Iliesiu:2024cnh].
@@ Appendix C Microcanonical approximations to length operators
Let $\Pi_{r}$ project onto a Hilbert space $\mathcal{H}_{r}$ spanned by $d$ consecutive energy levels from $E_{r}$ to $E_{r+d-1}$, each such set of levels constituting a microcanonical window, such that $\sum_{r}\Pi_{r}=1$. The index r=(m1)dr=(m-1)d represents the “base index” of the $m$-th microcanonical window in the energy spectrum. In practice, we will be interested in a subset of these microcanonical windows spanning some interval $E_{\min},E_{\max}]$; let $R$ be the set of values of the base index $r$ corresponding to this interval.
@@
Let $\mathcal{H}_{\text{mc}}(\nu_{\text{mc}})$ be the set of states that are supported on $\nu_{\text{mc}}$ consecutive microcanonical windows spanning $E_{\min},E_{\max}]$ and are unbiased in each window:
@@
$=\begin{dcases}\frac{\pi}{2},\ &\text{ for }n=m,\\ -\frac{\pi(1+(-1)^{m-n})}{d^{2}\sin^{2}\left(\frac{m-n}{d}\pi\right)},\ &\text{ for }n\neq m.\end{dcases}$ (C.13)
@@
In the second line, we have set q=mnq=m-n, used the evenness of the sum in $q$, and noted that $\operatorname{sinc}^{2}(0)=1$ by definition to allow $q=0$. Introducing the variable $\xi=q/d$, we can write the sum in the above expression as:
@@
for some small $\epsilon_{\beta}>0$ (note that the density of states is monotonically increasing with energy, so if this can be established for the $d$ lowest levels in $E_{\min},E_{\max}]$, we should expect it to be the case for [almost] all energy levels throughout the spectrum, especially given spectral rigidity, as successive sets of $d$ levels span increasingly narrower energy ranges). We get
@@
As we discuss in Sec. 4.2, most of the probability of the Hartle-Hawking state can be found in a finite ($\beta$-dependent) energy range, which we choose to be $E_{\min},E_{\max}]$, we can truncate the support of $\lvert\beta_{\text{mc}}\rangle$ to this range and normalize this truncated version so that it formally belongs to $\mathcal{H}_{\text{mc}}(\nu_{\mathrm{mc}})$. But it is also the case that to accurately capture expectation values such as $\langle x^{q}\rangle$, the choice of microcanonical windows that must be accounted for is rather different: see Appendix D.
@@ Appendix E Wormhole dynamics with Poisson statistics
At a technical level, we can simulate this system by following the procedure of Sec. 4.3 except with the CUE eigenphases $\varphi_{n}$ replaced by a uniform, random (and sorted) distribution of points in $0,2\pi)$. With Poisson statistics, we expect the initial state to spread out rapidly to the point of exhibiting no clear imprints of the classical dynamics near the Heisenberg time scale due to a lack of spectral rigidity, and the absence of a bump in $\langle x(t)\rangle$ (more precisely, the absence of an “equal” exploration of values below and above the saturation value) due to a lack of nearest-neighbor level repulsion. This is illustrated in Figs. 9 and 10, respectively for a microcanonical state and a Hartle-Hawking state.